Wheeler’s Bike Company manufactures custom racing bicycles. The company uses a job order cost system to determine the cost of each bike. Estimated costs and expenses for the coming year follow:

Bike parts $ 348,800
Factory machinery depreciation 53,000
Factory supervisor salaries 142,500
Factory direct labor 195,663
Factory supplies 44,900
Factory property tax 27,250
Advertising cost 27,500
Administrative salaries 46,000
Administrative-related depreciation 20,700
Total expected costs $ 906,313

Required:

1. Calculate the predetermined overhead rate per direct labor hour if the average direct labor rate is $11.31 per hour.

2. Determine the amount of applied overhead if 17,600 actual hours are worked in the upcoming year.
Answer #1

1.

Direct labor cost = $195,663

Direct labor rate = $11.31 per hour

Direct labor hours = Direct labor cost / Direct labor rate

= 195,663 / 11.31

= 17,300 hours

Estimated factory overheads = Factory machinery depreciation + Factory supervisor salaries + Factory supplies + Factory property tax

= 53,000+142,500+44,900+27,250

= $267,650

Predetermined overhead rate = Estimated factory overheads / Direct labor hours

= 267,650/17,300

= $15.47 per direct labor hour ( round off up to two decimals)

2.

Applied overhead = Actual direct labor hours x Predetermined overhead rate

= 17,600 x 15.47

= $272,272

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks